OpenStudy (anonymous):

how to solve the equation -1/x-4=2/x-5 please explain

OpenStudy (anonymous):

(-1/x)-4 = (2/x)-5 add both side by 4 and 1/x then we have 0=(3/x)-1 add 1 both side 1=3/x multiply x both side x=3